OpenStudy (anonymous):

solve for x: ln(5x-3)-ln(2x+1)=ln2

OpenStudy (anonymous):

the answer is 5, i'm just not sure on how to get there

OpenStudy (amistre64):

5x-3 ---- = 2 2x+1

OpenStudy (anonymous):

ln A - ln B = ln(A/B)

OpenStudy (amistre64):

5x-3 = 4x+2 x = 5